OpenStudy (anonymous):

2cos x^2 -4x^2 sin x^2=0 could someone please help me with this

OpenStudy (anonymous):

first you would want to make x^2 a variable say= y

OpenStudy (anonymous):

then you have a simpler equation

OpenStudy (anonymous):

i tried that too, im still stuck :(

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@AriPotta @Abhisar @BOOKER23 @DivineSolar @fatima0520 @hari5719 @shifuyanli

OpenStudy (bohotness):

   2cos²x + sinx - 2 = 0 2(1-sin²x) + sinx - 2 = 0  2 - 2sin²x + sinx - 2 = 0          -2sin²x + sinx = 0            2sin²x - sinx = 0         sinx(2sinx - 1) = 0          solutions are for: i. sinx = 0 ⇒ x = 0, π, 2π  ✔ ii. 2sinx - 1 = 0         2sinx = 1           sinx = 1/2               x = π/6  , 5π/6  ✔        ANSWER x = 0, π/6, 5π/6, π, 2π

OpenStudy (anonymous):

@bohotness I cant understand how you got:  2cos²x + sinx - 2 = 0

OpenStudy (anonymous):

@bohotness the question is actually: 2cos (x^2) -4x^2 sin (x^2)=0

OpenStudy (bohotness):

Okay

OpenStudy (loser66):

What are we supposed to do? solve for x?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

actually is this part of a longer problem but im stuck on that, i mean to look for the point of inflection of a graph and this is the second derivatice

OpenStudy (loser66):

oh man, please, post the original one.

OpenStudy (anonymous):

Sketch the graph of the function y=sin(x2) for −2π≤x≤2π

OpenStudy (loser66):

https://www.desmos.com/calculator/ln4cyvqy2e

OpenStudy (anonymous):

yes i know, but i have to show the work to get to the answer. thats why i posted this

OpenStudy (anonymous):

im stuck while trying to solve2cos (x^2) -4x^2 sin (x^2)=0

OpenStudy (loser66):

To me, just assign the value of x and calculate y. Only thing to do is finding out where y =0

OpenStudy (anonymous):

when you make it x , then equate it to both sides

OpenStudy (loser66):

The amplitude of the function is 1, hence the min/ max are just -1,1 sin (x^2)=0 iff x ^2 =0 or x^2 =pi , solve for them.

OpenStudy (anonymous):

then you can divide from there you may get tan

OpenStudy (loser66):

I think you make it more complex when you take first and second derivative to find max/min/ inflection one or concave up/down. No need to do that since trig functions have only one shape :) (the wave one)

OpenStudy (anonymous):

no iits not derivative if you do it that way you will get 2cosx=4xsinx

OpenStudy (anonymous):

then tan^-1((1/2(x))

OpenStudy (anonymous):

is equal to x

OpenStudy (anonymous):

sorry i think human calculator has done it all

OpenStudy (anonymous):

@Loser66 I am trying to find the exact coordinates of the point of inflection

OpenStudy (loser66):

How? look at this http://www.wolframalpha.com/input/?i=%28sin%28x^2%29%29%22%3D0